OpenStudy (anonymous):

Put the equation in the form ax^2+bx+c=0 (2x+1)^2-3x=5

OpenStudy (radar):

4x^2+x-4=0 Steps shown below.

OpenStudy (radar):

(2x+1)^2=4x^2 + 5x + 1 combinig with the additional terms: 4x^2+4x-3x+1-5=0 4x^2+x-4=0 did you follow with understanding?